# LogLLM: Log-based Anomaly Detection Using Large Language Models

### v4

@@ I Introduction
As we know, LLMs frequently face out-of-memory challenges due to their extensive parameter sizes [41]. Directly inputting the entire log sequence (by concatenating log messages into a long string) into Llama can lead to out-of-memory issues and potentially confuse the LLM, making it difficult to focus on key points for distinguishing anomalies. By adopting BERT to summarize each log message, LogLLM effectively mitigates these problems. Compared to other methods, LogLLM fully exploits the capabilities of LLMs for log-based anomaly detection. We conduct experiments across four public datasets, and the results demonstrate that LogLLM outperforms state-of-the-art methods. Even when handling unstable logs, where new log templates frequently emerge due to software evolution, it effectively captures the semantic meaning of log messages and detects anomalies accurately. The ablation study confirms the effectiveness of the three-stage training procedure.
@@ II-B LLMs for Log-based Anomaly Detection
Fine-tuning-based methods [32, 33, 34, 35, 3, 36, 37, 38, 39, 40] incorporate LLMs into deep neural networks and customize them to the user’s own dataset. Some methods [32, 33, 34, 35], although adopting transformer encoder-based LLMs for anomaly detection, do not capture the semantic information within log sequences. For example, LogBERT [32] and LAnoBERT [33] utilize BERT to reconstruct the input sequence of log template IDs (IDs of log string templates) and detect anomalies based on reconstruction errors, disregarding the semantic information. Other methods [3, 36, 37, 38, 39] use transformer encoder-based LLMs solely for extracting semantic information from log messages and then employ either smaller models [3, 36, 37, 38] or distance-based comparison [39] for classification. For instance, NeuralLog [3] leverages BERT to extract semantic vectors from raw log messages, which are subsequently used to detect anomalies via a transformer-based classification model. Similarly, RAPID [39] utilizes transformer encoder-based models to extract semantic vectors and performs anomaly detection by comparing each query log sequence with its nearest document log sequence. Hadadi et al. [40] directly input template sequences parsed from log sequences, into an LLM and fine-tune it to accurately predict sequence labels. However, this method is impractical because each template may be processed into multiple tokens by the LLM’s tokenizer, and a single template sequence can contain numerous log templates. Consequently, an excessive number of tokens can be generated for one template sequence, which LLMs often cannot process due to token (memory) limitations [41].
LogLLM is a fine-tuning-based method that utilizes BERT for extracting semantic vectors from log messages and Llama, a transformer decoder-based model, for log sequence classification. This method aligns the vector representation spaces of BERT and Llama using a projector. By adopting BERT, LogLLM effectively mitigates the out-of-memory issue caused by excessive tokens when directly tokenizing the entire log sequence with Llama’s tokenizer. Compared to other methods, LogLLM fully exploits the capabilities of LLMs for log-based anomaly detection.
@@ IV Methodology
In this section, we present our innovative anomaly detection framework, LogLLM. As illustrated in Fig. 3, the log sequence undergoes preprocessing using regular expressions before being fed into a deep neural network that integrates BERT [43], a projector, and Llama 3 [26] for log sequence classification. In the following sections, we will provide detailed insights into log sequence preprocessing, the architecture of the deep model, and the model training procedure.
@@ IV-B1 BERT
BERT generates a semantic vector by processing the semantic vector of the classification token ([CLS]) through a linear layer followed by a tanh activation function. Each log message, once preprocessed, is encoded into a semantic vector using the BERT tokenizer and BERT model. For a preprocessed log sequence, the output of BERT is a sequence of semantic vectors $C=(c_{1},c_{2},\dots,c_{N})\in\mathbb{R}^{N\times d_{BERT}}$, where $N$ represents the length of the log sequence (i.e., the number of log messages) and $d_{BERT}$ is the dimension of each semantic vector (i.e., hidden size). LogLLM utilizes the BERT base model [^1] https://huggingface.co/google-bert/bert-base-uncased, which consists of 12 layers of transformer encoders and 768 hidden units in each transformer. Therefore, $d_{BERT}$ is 768.
@@ IV-B3 Llama
The first component introduces the log sequence, such as "Below is a sequence of system log messages:". The second component comprises the token embeddings $E$ output by the projector. The third component queries whether the sequence is anomalous, asking, for instance, ". Is this sequence normal or anomalous?". The first and third components are fed into the Llama tokenizer and Llama embedding layer sequentially, producing $E_{1}\in\mathbb{R}^{A\times d_{Llama}}$ and $E_{3}\in\mathbb{R}^{Q\times d_{Llama}}$, where $A$ and $Q$ are the number of tokens produced by tokenizing the first and third components, respectively. Then, the token embeddings of the three components are concatenated, represented as $[E_{1}||E||E_{3}]\in\mathbb{R}^{(A+N+Q)\times d_{Llama}}$ and fed into Llama. We utilize Llama 3 8B model [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B, which consists of 32 transformer decoder-like layers and contains 8 billion parameters. Here, $d_{Llama}$ is 4096.
@@ V Experiments
  • RQ3: How does each stage of the three-stage training process influence the performance of LogLLM?
  • RQ4: How do different levels of minority class oversampling, determined by the hyperparameter $\beta$, affect the performance of LogLLM?
@@ V-B Experimental Settings
In our experiment, the hyperparameter $\beta$, which is described in Section IV-C1, is set to 30%. We use the AdamW optimizer [56] to train the model with a mini-batch size of 16. Unless otherwise specified, the training procedure is configured as follows: In the first stage, only 1,000 samples are involved with a learning rate of 5e-4. The second and third stages each consist of two epochs with a learning rate of 5e-5.
@@ V-E Performance Evaluation (RQ1)
TABLE II: Experimental results on HDFS, BGL, Liberty, and Thunderbird datasets.
@@
Effect of labeled anomalies (RQ2): As illustrated in Table II, in contrast to methods such as DeepLog, LogAnomaly, FastLogAD, LogBERT, and RAPID, which require clean datasets devoid of anomalies to build anomaly detection models, methods like PLELog, LogRobust, CNN, NeuralLog, and LogLLM demonstrate superior performance. These models are trained using not only normal samples but also labeled anomalies. For instance, these five methods achieve an average F1-score above 0.771 across four datasets, whereas others that do not utilize labeled anomalies perform poorly, with an average F1-score below 0.602 across four datasets. This demonstrates that incorporating labeled anomalies can provide a significant advantage to anomaly detection methods.
@@
V-F Effects of Different Preprocessing Techniques (RQ3)
TABLE IV: Effects of different preprocessing techniques on HDFS, BGL, Liberty, and Thunderbird datasets.
@@ V-F Effects of Different Preprocessing Techniques (RQ3)
V-G Ablation Study of the Training Procedure (RQ4)
TABLE V: Ablation study of the training procedure on HDFS, BGL, Liberty, and Thunderbird datasets.
@@ V-G Ablation Study of the Training Procedure (RQ4)
We investigate the effect of each training procedure through an ablation study. The results are presented in Table V, where ‘W/O’ denotes ‘without’. We have the following observations:

### v5

@@ I Introduction
As illustrated in Section V-G, LLMs frequently face out-of-memory challenges due to their extensive parameter sizes [41]. Directly inputting the entire log sequence (by concatenating log messages into a long string) into Llama can lead to out-of-memory issues and potentially confuse the LLM, making it difficult to focus on key points for distinguishing anomalies. By adopting BERT to summarize each log message, LogLLM effectively mitigates these problems. We conduct experiments across four public datasets, and the results demonstrate that LogLLM outperforms state-of-the-art methods. Even when handling unstable logs, where new log templates frequently emerge due to software evolution, it effectively captures the semantic meaning of log messages and detects anomalies accurately. The ablation study confirms the effectiveness of the three-stage training procedure.
@@ II-B LLMs for Log-based Anomaly Detection
Fine-tuning-based methods [32, 33, 34, 35, 3, 36, 37, 38, 39, 40] incorporate LLMs into deep neural networks and customize them to the user’s own dataset. Some methods [32, 33, 34, 35], although adopting transformer encoder-based LLMs for anomaly detection, do not capture the semantic information within log sequences. For example, LogBERT [32] and LAnoBERT [33] utilize BERT to reconstruct the input sequence of log template IDs (IDs of log string templates) and detect anomalies based on reconstruction errors, disregarding the semantic information. Other methods [3, 36, 37, 38, 39] use transformer encoder-based LLMs solely for extracting semantic information from log messages and then employ either smaller models [3, 36, 37, 38] or distance-based comparison [39] for classification. For instance, NeuralLog [3] leverages BERT to extract semantic vectors from raw log messages, which are subsequently used to detect anomalies via a transformer-based classification model. Similarly, RAPID [39] utilizes transformer encoder-based models to extract semantic vectors and performs anomaly detection by comparing each query log sequence with its nearest document log sequence. Hadadi et al. [40] directly input template sequences parsed from log sequences, into GPT models and fine-tune it to accurately predict sequence labels. However, this approach faces two key challenges. First, the boundaries between templates within the sequences are unclear, making it difficult for the model to learn the sequential dependencies. Second, each template may be tokenized into multiple tokens by the LLM’s tokenizer, and a single sequence can contain numerous log templates. As a result, an excessive number of tokens may be generated, often exceeding the token (memory) limits of LLMs [41], thereby restricting the length of sequences that can be processed. These two challenges are further demonstrated in Section V-G.
LogLLM is a fine-tuning-based method that utilizes BERT for extracting semantic vectors from log messages and Llama, a transformer decoder-based model, for log sequence classification. This method aligns the vector representation spaces of BERT and Llama using a projector. The use of BERT ensures clear boundaries between log messages, as each message is represented by a distinct embedding vector, thereby enhancing classification performance. Moreover, when memory and parameter size of Llama are held constant, this approach can handle longer sequences compared to directly tokenizing the entire log sequence using Llama’s tokenizer.
@@ IV Methodology
In this section, we present our innovative anomaly detection framework, LogLLM. As illustrated in Fig. 3, the log sequence undergoes preprocessing using regular expressions before being fed into a deep neural network that integrates BERT [43], a projector, and Llama [26] for log sequence classification. In the following sections, we will provide detailed insights into log sequence preprocessing, the architecture of the deep model, and the model training procedure.
@@ IV-B1 BERT
BERT generates a semantic vector by processing the semantic vector of the classification token ([CLS]) through a linear layer followed by a tanh activation function. Each log message, once preprocessed, is encoded into a semantic vector using the BERT tokenizer and BERT model. For a preprocessed log sequence, the output of BERT is a sequence of semantic vectors $C=(c_{1},c_{2},\dots,c_{N})\in\mathbb{R}^{N\times d_{BERT}}$, where $N$ represents the length of the log sequence (i.e., the number of log messages) and $d_{BERT}$ is the dimension of each semantic vector (i.e., hidden size).
@@ IV-B3 Llama
The first component introduces the log sequence, such as "Below is a sequence of system log messages:". The second component comprises the token embeddings $E$ output by the projector. The third component queries whether the sequence is anomalous, asking, for instance, ". Is this sequence normal or anomalous?". The first and third components are fed into the Llama tokenizer and Llama embedding layer sequentially, producing $E_{1}\in\mathbb{R}^{A\times d_{Llama}}$ and $E_{3}\in\mathbb{R}^{Q\times d_{Llama}}$, where $A$ and $Q$ are the number of tokens produced by tokenizing the first and third components, respectively. Then, the token embeddings of the three components are concatenated, represented as $[E_{1}||E||E_{3}]\in\mathbb{R}^{(A+N+Q)\times d_{Llama}}$ and fed into Llama.
@@ V Experiments
  • RQ3: How effective is the embedder for Llama?
  • RQ4: How does the size of the Llama model affect the performance of LogLLM?
  • RQ5: How does each stage of the three-stage training process influence the performance of LogLLM?
  • RQ6: How do different levels of minority class oversampling, determined by the hyperparameter $\beta$, affect the performance of LogLLM?
@@ V-B Experimental Settings
We conduct all experiments on a server equipped with an Intel Xeon Gold 6330 CPU (38 cores), 256GB of memory, and an NVIDIA A40 GPU with 48 GB of memory.
In our experiment, we utilize the BERT-base model [^1] https://huggingface.co/google-bert/bert-base-uncased and Llama-3-8B model [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as backbones. The hyperparameter $\beta$, which is described in Section IV-C1, is set to 30%. We use the AdamW optimizer [56] to train the model with a mini-batch size of 16. Unless otherwise specified, the training procedure is configured as follows: In the first stage, only 1,000 samples are involved with a learning rate of 5e-4. The second and third stages each consist of two epochs with a learning rate of 5e-5.
@@ V-E Performance Evaluation (RQ1)
TABLE II: Experimental results on HDFS, BGL, Liberty, and Thunderbird datasets. The best results are highlighted in bold.
@@
Effect of labeled anomalies: As illustrated in Table II, in contrast to methods such as DeepLog, LogAnomaly, FastLogAD, LogBERT, and RAPID, which require clean datasets devoid of anomalies to build anomaly detection models, methods like PLELog, LogRobust, CNN, NeuralLog, and LogLLM demonstrate superior performance. These models are trained using not only normal samples but also labeled anomalies. For instance, these five methods achieve an average F1-score above 0.771 across four datasets, whereas others that do not utilize labeled anomalies perform poorly, with an average F1-score below 0.602 across four datasets. This demonstrates that incorporating labeled anomalies can provide a significant advantage to anomaly detection methods.
@@
V-F Different Preprocessing Techniques (RQ2)
TABLE IV: Effects of different preprocessing techniques on HDFS, BGL, Liberty, and Thunderbird datasets. The best results are highlighted in bold.
@@ V-F Different Preprocessing Techniques (RQ2)
TABLE V: Effects of the embedder (BERT & adapter) and LLaMA model size, where ‘Mem.’ indicates GPU memory usage (GB), and ‘Tim.’ indicates training time (Minutes). ‘-’ indicates an out-of-memory (OOM) error.
HDFS BGL Liberty Thunderbird
Prec. Rec. F1 Mem. Tim. Prec. Rec. F1 Mem. Tim. Prec. Rec. F1 Mem. Tim. Prec. Rec. F1 Mem. Tim.
L.-1B 0.986 0.995 0.991 16.5 1022.1 - - - - - 0.960 0.699 0.809 42.6 443.2 1.000 0.724 0.840 44.5 1732.1
Emb. & L.-1B 0.996 0.996 0.996 8.0 1412.2 0.734 0.944 0.825 32.4 187.1 0.950 0.905 0.927 29.3 173.2 0.875 0.966 0.918 32.4 715.1
L.-8B 0.988 0.997 0.992 43.0 4712.1 - - - - - - - - - - - - - - -
Emb. & L.-8B 0.994 1.000 0.997 16.6 2168.2 0.861 0.979 0.916 38.0 396.2 0.992 0.926 0.958 36.1 412.1 0.966 0.966 0.966 38.2 1284.2
@@
V-G Effect of the Embedder (RQ3)
We investigate whether the embedder (BERT and adapter) is necessary for LogLLM. The results are presented in Table V. ‘L.-1B’ refers to directly inputting the log sequence (by concatenating log messages with semicolons (;) as separators into a long string) into the ‘Llama-3.2-1B’ model  [^3] https://huggingface.co/meta-llama/Llama-3.2-1B. ‘Emb. & L.-1B’ represents LogLLM based on ‘Llama-3.2-1B’.
As expected, with the assistance of the embedder, the model requires less GPU memory, thereby avoiding out-of-memory (OOM) errors. Additionally, it enhances model performance by clarifying the boundaries between messages within a sequence. This improved representation enables the LLM to capture sequential dependencies better.
V-H Effect of the Llama Model Size (RQ4)
As shown in Table V, larger LLaMA model sizes lead to better performance, at the cost of increased GPU memory usage and longer training times.
On average, compared to using Llama-3.2-1B, adopting Llama-3-8B improves the F1-score by 4.3%, but increases GPU memory usage by 7.7 GB and extends training time by 443.2 minutes.
V-I Ablation Study of the Training Procedure (RQ5)
TABLE VI: Ablation study of the training procedure on HDFS, BGL, Liberty, and Thunderbird datasets. The best results are highlighted in bold.
@@ V-I Ablation Study of the Training Procedure (RQ5)
We investigate the effect of each training procedure through an ablation study. The results are presented in Table VI, where ‘W/O’ denotes ‘without’. We have the following observations: